Say you’re the owner of a car dealership and you want to find out the valuation that Kelley Blue Book uses as new models of cars are introduced Find the depreciation deduction, using the declining balance (DB) method with a 200% DB ratio, associated with the fourth year for a car that cost $41,000 and has an estimated MV of $8,300 at the end of its eight-year useful life. Express your answer in terms of dollars and cents, rounded to the nearest cent (e.g., 1234.56).
Answer #1

As it has a 200% DB Ratio, so the depreciation method to be used is "Double Declining Method". According to this method:

Depreciation in year x = (2/depreciable life in years)*Book value at the beginning of year x

Here, Book Value at the beginning = $ 41,000

Depreciable life = 8 years

Depreciated value for each year:

Year 1: (2/8) * 41000 = $ 10,250

Year 2: (2/7) * (41000 - 10250) = $ 8,785

Year 3: (2/6) * (41000 - 10250 - 8785) = $ 7,322

Since, the car is in its 4th year, the depreciation for the 4th year will be ignored.

So the valuation of the car at present is= 41000 - (10250 + 8785 + 7322) = $ 14,643